# A Survey of Self-Evolving Agents: On Path to Artificial Super Intelligence

### v3

@@ 7.1 Evaluation Goals and Metrics
While Adaptivity and Retention focus on in-domain performance, Generalization is a pivotal measure of a self-evolving agent’s ability to apply its accumulated knowledge to new, unseen domains or tasks. A truly intelligent agent should not only perform well within its familiar territory but also demonstrate a capacity for cross-domain generalization. This capability can be evaluated by assessing an agent’s performance on a diverse set of tasks that span multiple task distributions and domains. Common approaches include computing aggregate performance metrics (e.g., mean success rates) across multi-domain test suites liu2023agentbench , sun2023adaplanner , and conducting out-of-domain evaluations using held-out task distributions that simulate real-world novelty scenarios hu2024agentgen ; peng2025dlpo .
Efficiency
Efficiency quantifies the resourcefulness of a self-evolving agent. As agents operate continuously and make decisions autonomously, it is essential to evaluate the cost and speed of their evolutionary process. These metrics are particularly important for practical, real-world applications where resources like computation and time are finite. Key indicators include token consumption hu2024treeplannerefficientcloselooptask , which measures the computational cost of an agent’s reasoning and generation steps), time consumption lu2024axis , the number of steps wang2023voyager and the number of interaction with the tools (e.g. tool productivity)  (wang2025otcpo, ) required to complete a task, which rewards agents for completing tasks in the fewest possible resource consumption. Several key metrics quantify agent efficiency in task execution, including token consumption (measuring computational overhead in reasoning and generation) hu2024treeplannerefficientcloselooptask , time expenditure lu2024axis , the number of required stepswang2023voyager , and tool interaction frequency (wang2025otcpo, ). These indicators collectively assess an agent’s ability to minimize resource utilization while maintaining task performance, with lower values generally reflecting more efficient operation.
Safety
@@
From the perspective of self-evolving, the Safety domain critically examines whether these agents develop unsafe or undesirable behavioral patterns throughout their continuous evolution. This dimension assesses an agent’s adherence to predefined rules and its propensity for harmful actions. Key metrics in evaluating safety of self-evolving agents may include: (1) Safety Score zhang2024agentsafety , measures the proportion of test cases where the agent’s behavior is labeled “safe” ; (2) Harm Score andriushchenko2024agentharm , computes via a detailed manually written grading rubric where outputs earn partial credit whenever some but not all harmful criteria are triggered; (3) Completion Under Policy (CuP) levy2024st , assesses whether an agent successfully completes a task while strictly adhering to a given set of rules or policies ; (4) Risk Ratio levy2024st , calculates the frequency of an agent’s rule violations along a specific dimension, providing a quantitative measure of non-compliance ; (5) Refusal Rate zhang2024agentasb ; andriushchenko2024agentharm , evaluates the proportion of tasks an agent refuses to perform due to their aggressive, malicious, or otherwise unsafe nature; (6) Leakage Rate shao2024privacylens , tracks how often an agent unintentionally leaks sensitive or private information.
@@
Adaptivity Success Rate by Iteration Steps hu2024automated ; wang2024agent ; zheng2025lifelongagentbench Performance in downstream tasks as a function of the agent’s interaction history
@@
Adaptation Speed wang2023voyager How quickly an agent reaches a certain performance threshold or converges to an optimal strategy within a given adaptation period
@@
Retention Forgetting (FGT) zheng2025lifelong The average accuracy drop on old tasks after an agent learns a new one, measuring whether useful experience is successfully maintained
@@
Backward Transfer (BWT) zheng2025lifelong The average accuracy improvement on old tasks due to the experience gained from new tasks
@@
Generalization Aggregate Performance liu2023agentbench ; sun2023adaplanner Mean success rates or other performance indicators across multi-domain test suites to gauge overall proficiency
Out-of-Domain (OOD) Performance hu2024agentgen ; peng2025dlpo The agent’s performance in held-out task distributions
@@
Efficiency Token Consumption hu2024treeplannerefficientcloselooptask Computational overhead in reasoning and generation steps
Time Expenditure lu2024axis Total duration required for task completion
@@
Number of Steps wang2023voyager Minimal actions needed to accomplish objectives
@@
Tool Productivity wang2025otcpo The ratio between task benefit (e.g., answer accuracy) and tool usage cost (e.g., number of tool calls)
@@
Safety Safety Score zhang2024agentsafety Proportion of test cases where agent behavior meets predefined safety criteria
@@
Harm Score andriushchenko2024agentharm Graded assessment of harmful outputs based on violation severity
@@
Completion Under Policy (CuP) levy2024st Task success rate while complying with specified constraints
@@
Risk Ratio levy2024st Frequency of policy violations per interaction opportunity
@@
Refusal Rate zhang2024agentasb ; andriushchenko2024agentharm Percentage of tasks declined due to safety concerns
@@
Leakage Rate shao2024privacylens Incidence of unintended sensitive information disclosure
@@ 7.2 Evaluation Paradigm
The evaluation of self-evolving agents, given their continuous learning paradigm, necessitates a multi-faceted approach that extends beyond traditional static assessments. Current evaluation paradigm can be broadly categorized based on the temporal scope of the assessment: Static Assessment, Short-horizon Adaptive Assessment, and Long-horizon Lifelong Learning Ability Assessment. Each category addresses different aspects of an agent’s evolving capabilities, from its instantaneous performance to its long-term learning trajectory.
@@ 7.2.1 Static Assessment
This category assesses an agent’s end-to-end proficiency in completing tasks across various real-world or simulated environments. In scientific data analysis and machine learning engineering, benchmarks like ScienceAgentBenchchen2024scienceagentbench and MLE-Benchchan2024mle test agents’ ability to generate and execute code for data analysis and solve Kaggle-style problems. For web search/Browse, environments such as WebShopyao2022webshop , WebArenazhou2023webarena , X-WebAgentBench wang2025x , Mind2Webdeng2023mind2web , and BrowseCompwei2025browsecomp simulate realistic web interactions, complex Browse scenarios, and task completion under security constraints. In software engineering, the SWE-bench seriesjimenez2023swe ; swebenchverified ; aleithan2024swe ; yang2024swe uses real GitHub issues to assess agents’ code repair capabilities. For computer-use interactions, OSWorldxie2024osworld offers a unified environment for open-ended tasks involving various desktop and web applications. Specialized domains like marketing also feature benchmarks such as xbenchchen2025xbench . Beyond specific domains, generalist agent benchmarks like AgentBenchliu2023agentbench , GAIAmialon2023gaia , and TheAgentCompanyxu2024theagentcompany evaluate broad problem-solving abilities across multiple knowledge domains and professional tasks, simulating real-world demands on general AI assistants.
@@
Beyond end-to-end task completion, assessing an agent’s underlying core competencies is crucial. These benchmarks evaluate fundamental capabilities that contribute to an agent’s overall intelligence and self-evolutionary potential. As for Planning, Benchmarks such as PlanBenchvalmeekam2023planbench , Natural Planzheng2024natural , AutoPlanBenchstein2025automating , and ACPBenchkokel2025acpbench comprehensively evaluate an agent’s ability to understand dynamic environments, devise strategies, decompose complex problems, and execute reasoning in various planning domains. For Tool Usage, simple benchmarks like ToolAlpacatang2023toolalpaca and ToolBenchqin2023toolllm test basic selection and parameter mapping, while more complex ones like ToolSandboxlu2024toolsandbox , Seal-Toolswu2024seal , API-Bankli2023api , T-Evalchen2023t , $\tau$-Benchyao2024tau , AceBenchchen2025acebench ) simulate real-world scenarios involving multi-turn interactions, implicit state dependencies, and nested calls. Memory Management benchmarks such as LTMbenchmarkcastillo-bolado2024beyond , MemoryAgentBenchhu2025evaluating , and StoryBenchwan2025storybench evaluate the agent’s capacity to retain and utilize information across multi-turn interactions, dynamic scenarios, and long-range dependencies. For evaluating Multi-Agent Collaboration, benchmarks such as MultiAgentBenchzhu2025multiagentbench and SwarmBenchruan2025benchmarking assess coordination, communication, and emergent swarm intelligence in both collaborative and competitive settings.
Metrics for Static Assessment
@@
ScienceAgentBench chen2024scienceagentbench Scientific Data Analysis Adaptivity, Efficiency Valid Execution Rate, Success Rate, CodeBERTScore, API Cost 102 Static
MLE-Bench chan2024mle ML-Engineering Adaptivity - 75 Static
DS-Bench jing2025dsbenchfardatascience Data Science Adaptivity Task Success Rate, Cost, Inference Time, Competition-level Accuracy 540 Static
SWE-bench jimenez2023swe Software Engineering Adaptivity Pass Rate 2,294 Static
OSWorld xie2024osworld Computer-Use / GUI Adaptivity Success Rate 369 Static
Mobile-Eval-E wang2025mobile Computer-Use / GUI Adaptivity, Efficiency Action Accuracy, Reflection Accuracy, Termination Error 25 Static, Short-horizon
WebShop yao2022webshop Web Search / Browse Adaptivity Success Rate 12,087 Static
WebArena zhou2023webarena Web Search / Browse Adaptivity Success Rate 812 Static
WebWalkerQA wu2025webwalker Web Search / Browse Adaptivity, Efficiency Accuracy, Action Count 680 Static
ST-WebAgentBench levy2024st Web Search / Browse Safety Completion under Policy 235 Static
xbench chen2025xbench Web Search / Browse Adaptivity LLM-Judge Score 100 Static
BrowseComp wei2025browsecomp Web Search / Browse Adaptivity Accuracy 1,266 Static
Agent-SafetyBench zhang2024agentsafety General Safety Safety Sore 20,000 Static
LifelongAgentBench zheng2025lifelongagentbench General Adaptivity, Retention, Generalization Success Rate 1396 Long-horizon
AgentBench liu2023agentbench General Adaptivity, Generalization Success Rate, F1, Reward, Game Progress 1360 Static
GAIA mialon2023gaia General Adaptivity Accuracy 466 Static
TheAgentCompany xu2024theagentcompany General Adaptivity, Efficiency Completion Score, Steps, Cost paer Instance 175 Static
EvaLearn dou2025evalearnquantifyinglearningcapability General Adaptivity, Efficiency Accuracy, Slope, Position of 1st solution, Num of consecutive solutions 648 Long-Horizon
PlanBench valmeekam2023planbench Planning Adaptivity Accuracy $\sim$26,250 Static, Short-horizon
Natural Plan zheng2024natural Planning Adaptivity Exact Match 3,600 Static
ACPBench kokel2025acpbench Planning Adaptivity, Generalization Accuracy 3,720 Static
AppBench (wang-etal-2024-appbench, ) Planning Adaptivity Success Rate, F1 800 Static
ToolBench qin2023toolllm Tool Usage Adaptivity Pass Rate, Win Rate 126,486 Static
ToolSandbox lu2024toolsandbox Tool Usage Adaptivity Similarity Score 1,032 Static
Seal-Tools wu2024seal Tool Usage Adaptivity Accuracy, P/R/F1 14,076 Static
API-Bank li2023api Tool Usage Adaptivity Accuracy, Rouge 4,125 Static
T-Eval chen2023t Tool Usage Adaptivity Domain-Specific Score 23,305 Static
$\tau$-Bench yao2024tau Tool Usage Adaptivity Pass^k 165 Static
AceBench chen2025acebench Tool Usage Adaptivity Accuracy 2,000 Static
LTMBenchmark castillo-bolado2024beyond Agent Memory Retention, Efficiency Score, Accuracy, GoodAI LTM Score, Speed, Cost, Verbosity 30 Long-Horizon
StoryBench wan2025storybench Agent Memory Retention, Efficiency Accuracy, First-Try Accuracy, Longest Corr, Retry Count, Runtime Cost, Token Cons, 311 scene nodes, 86 choice nodes Short-Horizon, Long-Horizon
MemoryAgentBench hu2025evaluating Agent Memory Adaptivity SubEM, Recall, ROUGE F1, Accuracy, Recall@5, Model Based Acc/F1 2200 Static, Short-horizon
MultiAgentBench zhu2025multiagentbench Multi-Agent Collaboration Adaptivity KPI, Text-Based Score, Communication Score, Planing Score, Coordination Score 100 Static
SwarmBench ruan2025benchmarking Multi-Agent Collaboration Adaptivity Perspective-specific Metrics 5 Short-horizon

### v4

@@ 7.1.3 Generalization
While Adaptivity and Retention focus on in-domain performance, Generalization is a pivotal measure of a self-evolving agent’s ability to apply its accumulated knowledge to new, unseen domains or tasks. A truly intelligent agent should not only perform well within its familiar territory but also demonstrate a capacity for cross-domain generalization. This capability can be evaluated by assessing an agent’s performance on a diverse set of tasks that span multiple task distributions and domains. Common approaches include computing aggregate performance metrics (e.g., mean success rates) across multi-domain test suites (Liu et al., 2023b; Sun et al., 2023), and conducting out-of-domain evaluations using held-out task distributions that simulate real-world novelty scenarios (Hu et al., 2024b; Peng et al., 2025).
Coverage Gaps
Generalization is typically evaluated through multi-domain test suites (AgentBench (Liu et al., 2023b), TheAgentCompany (Xu et al., 2024b) and held-out task distributions, measuring an agent’s ability to transfer knowledge across domains. Current evaluations, however, rely on static snapshots: agents are tested once on diverse tasks without tracking whether cross-domain transfer degrades as they evolve over extended learning trajectories. Technological progress further reduces the discriminative power of older tasks, as modern agents benefit from algorithmic advances unavailable to earlier systems (Chan et al., 2024), while training data contamination complicates cross-domain assessment (Chan et al., 2024). No existing benchmark examines whether agents preserve generalization breadth as they specialize within domains, or whether knowledge acquired in one domain continues to transfer after hundreds of learning episodes.
7.1.4 Efficiency
Goals & Metrics
Efficiency quantifies the resourcefulness of a self-evolving agent during its learning and operation. As agents operate continuously and make decisions autonomously, it is essential to evaluate the cost and speed of their evolutionary process.
These metrics in Table 5 are particularly important for practical, real-world applications where resources like computation, memory, time, and human effort are finite:
  • Token consumption (Hu et al., 2024a): Total tokens used in reasoning, generation, and memory operations
  • Time consumption (Lu et al., 2024b): Wall-clock time required to complete tasks or reach performance thresholds
  • Step count (Wang et al., 2023a): Number of interaction rounds needed for task completion
  • Tool calls: Number of external API/environment invocations, can be combined with performance gains to assess Tool Productivity (TP) (Wang et al., 2025g), formulated as:
    $\quad\mathrm{TP}_{\mathdollar}=\frac{\Delta\text{score}}{\sum_{i}\text{cost}(\text{tool}_{i})}.$
  • Memory growth: Expansion of persistent memory and context window over the agent’s operational duration
  • Human oversight: Time and effort required for human intervention, including review, labeling, and corrective guidance
Table 5: Refined cost taxonomy and units for self-evolving agents, with Real-world example (Fan et al., 2025): On SWE-bench (Jimenez et al., 2024), SWE-Agent (Yang et al., 2024) + Qwen3-32B (440K tokens, 35 calls, 28% success) vs. GPT-4o-mini (8.1M tokens, 181 calls, 10% success)—model-scaffold synergy critical for efficiency.
Dimension Symbol Typical Unit Example Measurement Worked Example
Token usage $C_{\text{token}}$ #tokens Prompt + completion tokens per episode/step 440K vs. 8.1M tokens per task; 18× difference
Step count / latency $C_{\text{step}}$ #turns / rounds Reasoning/interaction turns 35 vs. 181 API calls per task; Quality vs. quantity trade-off
Wall-clock time $C_{\text{time}}$ #seconds End-to-end elapsed runtime incl. I/O 12 vs. 200+ turns per task; Simple vs. complex bugs
Tool/API calls $C_{\text{tool}}$ #calls External function/environment calls 15 vs. 38 calls (AutoCodeRover); High-quality reasoning reduces iterations
Memory growth $C_{\text{mem}}$ #tokens / MB Persistent memory; context window expansion Linear growth per call; Token snowball effect amplifies invalid context
Human oversight $C_{\text{human}}$ #hours / tokens Review, labeling, red-team; guidance tokens Failed attempts: 4× cost vs. successful ones; Expensive failures pattern
To relate efficiency to performance gains, we report Cost-per-Gain (CPG) as:
$\mathrm{CPG}_{t}=\frac{\text{Total Cost}_{t}}{\text{Performance Gain}_{t}+\epsilon},$
where cost can be measured in tokens, time, memory, human effort, or a normalized composite (e.g., monetary cost), and performance gain is the improvement over baseline at horizon $t$. Lower CPG indicates more efficient learning.
Coverage Gaps
Efficiency suffers from sparse and inconsistent reporting of evolution costs. While benchmark papers occasionally document aggregate resource consumption during evaluation (Chan et al., 2024), they rarely decompose costs into evolution-specific components: tokens consumed during self-reflection or experience replay, wall-clock time spent on architecture search or memory updates, tool invocations triggered by autonomous exploration. Cost constraints in human baseline collection (Xu et al., 2024b) reflect broader accessibility challenges but do not illuminate the efficiency of the evolution process itself. More fundamentally, standard evaluation protocols permit unconstrained optimization—agents maximize task success without facing the hard token budgets, iteration limits, or latency constraints that govern real-world deployment.
7.1.5 Safety
Goals & Metrics
@@ 7.1.5 Safety
From the perspective of self-evolving, the Safety domain critically examines whether these agents develop unsafe or undesirable behavioral patterns throughout their continuous evolution. This dimension assesses an agent’s adherence to predefined rules and its propensity for harmful actions. Key metrics in evaluating safety of self-evolving agents may include: (1) Safety Score (Zhang et al., 2024i), which measures the proportion of test cases where the agent’s behavior is labeled “safe”; (2) Harm Score (Andriushchenko et al., 2024), computed via a detailed manually written grading rubric where outputs earn partial credit whenever some but not all harmful criteria are triggered; (3) Completion Under Policy (CuP) (Levy et al., 2024), which assesses whether an agent successfully completes a task while strictly adhering to a given set of rules or policies; (4) Risk Ratio (Levy et al., 2024), which calculates the frequency of an agent’s rule violations along a specific dimension, providing a quantitative measure of non-compliance; (5) Refusal Rate (Zhang et al., 2024b; Andriushchenko et al., 2024), which evaluates the proportion of tasks an agent refuses to perform due to their aggressive, malicious, or otherwise unsafe nature; (6) Leakage Rate (Shao et al., 2024a), which tracks how often an agent unintentionally leaks sensitive or private information.
Coverage Gaps
Safety evaluation predominantly captures risks in isolated episodes. Agent-SafetyBench identifies two core deficiencies: inadequate robustness when deploying tools across varied contexts, and limited recognition of potential hazards in specific operational scenarios (Zhang et al., 2024i). In multi-agent settings, SwarmBench reveals fundamental coordination challenges where local interactions fail to produce coherent collective strategies, with agents unable to maintain shared situational understanding necessary for safe collaborative operation (Ruan et al., 2025). Yet no benchmark tracks safety trajectories over extended evolution—whether risks accumulate through repeated exposure to edge cases, or whether unsafe behaviors emerge through autonomous exploration and self-directed learning.
@@
Adaptivity Success Rate by Iteration Steps (Hu et al., 2024c; Wang et al., 2024j; Zheng et al., 2025b) Performance in downstream tasks as a function of the agent’s interaction history
@@
Adaptation Speed (Wang et al., 2023a) How quickly an agent reaches a certain performance threshold or converges to an optimal strategy within a given adaptation period
@@
Retention Forgetting (FGT) (Zheng et al., 2025c) The average accuracy drop on old tasks after an agent learns a new one, measuring whether useful experience is successfully maintained
@@
Backward Transfer (BWT) (Zheng et al., 2025c) The average accuracy improvement on old tasks due to the experience gained from new tasks
@@
Generalization Aggregate Performance (Liu et al., 2023b; Sun et al., 2023) Mean success rates or other performance indicators across multi-domain test suites to gauge overall proficiency
Out-of-Domain (OOD) Performance (Hu et al., 2024b; Peng et al., 2025) The agent’s performance in held-out task distributions
@@
Efficiency Token Consumption (Hu et al., 2024a) Computational overhead in reasoning and generation steps
Time Expenditure (Lu et al., 2024b) Total duration required for task completion
@@
Number of Steps (Wang et al., 2023a) Minimal actions needed to accomplish objectives
@@
Tool Productivity (Wang et al., 2025g) The ratio between task benefit (e.g., answer accuracy) and tool usage cost (e.g., number of tool calls)
@@
Safety Safety Score (Zhang et al., 2024i) Proportion of test cases where agent behavior meets predefined safety criteria
@@
Harm Score (Andriushchenko et al., 2024) Graded assessment of harmful outputs based on violation severity
@@
Completion Under Policy (CuP) (Levy et al., 2024) Task success rate while complying with specified constraints
@@
Risk Ratio (Levy et al., 2024) Frequency of policy violations per interaction opportunity
@@
Refusal Rate (Zhang et al., 2024b; Andriushchenko et al., 2024) Percentage of tasks declined due to safety concerns
@@
Leakage Rate (Shao et al., 2024a) Incidence of unintended sensitive information disclosure
ScienceAgentBench(Chen et al., 2024e) Scientific Data Analysis Adaptivity, Efficiency Valid Execution Rate, Success Rate, CodeBERTScore, API Cost 102 Static
MLE-Bench(Chan et al., 2024) ML-Engineering Adaptivity 75 Static
DS-Bench(Jing et al., 2025) Data Science Adaptivity Task Success Rate, Cost, Inference Time, Competition-level Accuracy 540 Static
SWE-bench(Jimenez et al., 2023) Software Engineering Adaptivity Pass Rate 2,294 Static
OSWorld(Xie et al., 2024) Computer-Use / GUI Adaptivity Success Rate 369 Static
Mobile-Eval-E(Wang et al., 2025n) Computer-Use / GUI Adaptivity, Efficiency Action Accuracy, Reflection Accuracy, Termination Error 25 Static, Short-horizon
WebShop(Yao et al., 2022) Web Search / Browse Adaptivity Success Rate 12,087 Static
WebArena(Zhou et al., 2023) Web Search / Browse Adaptivity Success Rate 812 Static
WebWalkerQA(Wu et al., 2025) Web Search / Browse Adaptivity, Efficiency Accuracy, Action Count 680 Static
ST-WebAgentBench(Levy et al., 2024) Web Search / Browse Safety Completion under Policy 235 Static
xbench(Chen et al., 2025f) Web Search / Browse Adaptivity LLM-Judge Score 100 Static
BrowseComp(Wei et al., 2025b) Web Search / Browse Adaptivity Accuracy 1,266 Static
Agent-SafetyBench(Zhang et al., 2024i) General Safety Safety Score 20,000 Static
LifelongAgentBench(Zheng et al., 2025b) General Adaptivity, Retention, Generalization Success Rate 1396 Long-horizon
AgentBench(Liu et al., 2023b) General Adaptivity, Generalization Success Rate, F1, Reward, Game Progress 1360 Static
GAIA(Mialon et al., 2023) General Adaptivity Accuracy 466 Static
TheAgentCompany(Xu et al., 2024b) General Adaptivity, Efficiency Completion Score, Steps, Cost per Instance 175 Static
EvaLearn(Dou et al., 2025) General Adaptivity, Efficiency Accuracy, Slope, Position of 1st solution, Num of consecutive solutions 648 Long-horizon
PlanBench(Valmeekam et al., 2023) Planning Adaptivity Accuracy $\sim$26,250 Static, Short-horizon
Natural Plan(Zheng et al., 2024) Planning Adaptivity Exact Match 3,600 Static
ACPBench(Kokel et al., 2025) Planning Adaptivity, Generalization Accuracy 3,720 Static
AppBench (Wang et al., 2024b) Planning Adaptivity Success Rate, F1 800 Static
ToolBench(Qin et al., 2023) Tool Usage Adaptivity Pass Rate, Win Rate 126,486 Static
ToolSandbox(Lu et al., 2024a) Tool Usage Adaptivity Similarity Score 1,032 Static
Seal-Tools(Wu et al., 2024) Tool Usage Adaptivity Accuracy, P/R/F1 14,076 Static
API-Bank(Li et al., 2023) Tool Usage Adaptivity Accuracy, ROUGE 4,125 Static
T-Eval(Chen et al., 2023) Tool Usage Adaptivity Domain-Specific Score 23,305 Static
$\tau$-Bench(Yao et al., 2024) Tool Usage Adaptivity Pass^k 165 Static
AceBench(Chen et al., 2025a) Tool Usage Adaptivity Accuracy 2,000 Static
LTMBenchmark(Castillo-Bolado et al., 2024) Agent Memory Retention, Efficiency Score, Accuracy, GoodAI LTM Score, Speed, Cost, Verbosity 30 Long-horizon
StoryBench(Wan and Ma, 2025) Agent Memory Retention, Efficiency Accuracy, First-Try Accuracy, Longest Corr, Retry Count, Runtime Cost, Token Consumption 311 scene nodes, 86 choice nodes Short-horizon, Long-horizon
MemoryAgentBench(Hu et al., 2025) Agent Memory Adaptivity SubEM, Recall, ROUGE F1, Accuracy, Recall@5, Model-based Acc/F1 2200 Static, Short-horizon
MultiAgentBench(Zhu et al., 2025) Multi-Agent Collaboration Adaptivity KPI, Text-Based Score, Communication Score, Planning Score, Coordination Score 100 Static
SwarmBench(Ruan et al., 2025) Multi-Agent Collaboration Adaptivity Perspective-specific Metrics 5 Short-horizon
Table 8: Goal-to-benchmark mapping with limitations and coverage gaps.
Goal Benchmarks Limitations Coverage Gaps
Adaptivity SWE-bench, WebArena, MLE-Bench, ScienceAgentBench, GAIA Restricted programming languages and execution time limits (Chen et al., 2024e); Simplified problem specifications vs. authentic R&D ambiguity (Chan et al., 2024); Weak sustained reasoning and decision-making (Liu et al., 2023b) Open-ended exploration without predetermined objectives; Long-horizon adaptation under non-stationary distributions
Retention LifelongAgentBench, LTMBenchmark, MemoryAgentBench Persistent challenges in dynamic memory and long-range consistency (Hu et al., 2025); Replay buffer trade-offs with context overflow (Zheng et al., 2025b); Limited robustness testing due to costs (Castillo-Bolado et al., 2024) Episodic designs reset state between tasks; No retention with safety constraints
Generalization AgentBench, GAIA, TheAgentCompany, MLE-Bench Progressive task obsolescence from algorithmic advances (Chan et al., 2024); Training data contamination risks (Chan et al., 2024) Temporal robustness under distribution drift; Adversarial co-evolutionary assessment
Efficiency MLE-Bench, TheAgentCompany High resource demands limit accessibility (Chan et al., 2024); Cost constraints prevent baseline collection (Xu et al., 2024b) No enforced budgets during evolution; Multi-objective constraints absent
Safety Agent-SafetyBench, SwarmBench Inadequate tool robustness and risk awareness (Zhang et al., 2024i); Local-global coordination disconnect (Ruan et al., 2025) Static evaluation only; No long-horizon safety drift tracking; Co-evolutionary safety unexplored
7.1.6 Self-Directedness and Evaluation Trade-offs
A defining characteristic of self-evolving agents is their capacity for autonomous exploration and self-directed evolution, distinguishing them from passive learning paradigms. The degree of self-directedness, specifically whether the agent autonomously generates tasks and evolution strategies or follows externally provided curricula, creates fundamental trade-offs across evaluation dimensions.
Highly self-directed systems demonstrate substantial performance gains through autonomous curriculum generation. WebRL improved from 4.8% to 42.4% by self-generating tasks from exploration failures (Qi et al., 2024), while SEAgent achieved 23.2 percentage point improvements, advancing from 11.3% to 34.5% via autonomous software exploration and curriculum evolution (Sun et al., 2025c). However, autonomous evolution incurs measurable risks. Alignment faking rates escalated from 12% to 78% when agents autonomously evolved under conflicting objectives (Greenblatt et al., 2024), illustrating safety challenges when evolution proceeds with minimal external oversight.
Despite these documented impacts, the field lacks standardized metrics for quantifying self-directedness in evolution. To enable fair comparison, we recommend transparently reporting three aspects of the evolution process. (1) specify whether evolution strategies and task sequences are predetermined, procedurally sampled, or autonomously generated by the agent; (2) Document the source of feedback signals, indicating whether they come from external human labels, rule-based evaluators, or self-generated reflection; (3) Report the frequency of external interventions in the evolution process. Clearly documenting the degree of autonomous control is essential for interpreting performance claims, as gains may reflect true self-evolution capability or extensive external guidance.
Such trajectory-centric evaluation aligns with the core objective of assessing self-evolving agents: measuring not just what they achieve, but how autonomously they learn to evolve.
@@ 7.2 Evaluation Paradigm
The evaluation of self-evolving agents, given their continuous learning paradigm, necessitates a multi-faceted approach that extends beyond traditional static assessments. Current evaluation paradigm can be broadly categorized based on the temporal scope of the assessment: Static Assessment, Short-horizon Adaptive Assessment, and Long-horizon Lifelong Learning Ability Assessment. Each category addresses different aspects of an agent’s evolving capabilities, from its instantaneous performance to its long-term learning trajectory. Terminology note. Here, “continuous” refers to system-level self‑evolution (changes in policies, memories, skills, tools, or procedures across episodes), not necessarily model‑parameter continual learning. Our evaluation therefore traces trajectories across interaction episodes and evolving environments, independent of whether weights are updated.
@@ 7.2.1 Static Assessment
This category assesses an agent’s end-to-end proficiency in completing tasks across various real-world or simulated environments. In scientific data analysis and machine learning engineering, benchmarks like ScienceAgentBench (Chen et al., 2024e) and MLE-Bench (Chan et al., 2024) test agents’ ability to generate and execute code for data analysis and solve Kaggle-style problems. For web search/browsing, environments such as WebShop (Yao et al., 2022), WebArena (Zhou et al., 2023), X-WebAgentBench (Wang et al., 2025i), Mind2Web (Deng et al., 2023), and BrowseComp (Wei et al., 2025b) simulate realistic web interactions, complex browsing scenarios, and task completion under security constraints. In software engineering, the SWE-bench series (Jimenez et al., 2023; Openai, 2024; Aleithan et al., 2024; Yang et al., 2024) uses real GitHub issues to assess agents’ code repair capabilities. For computer-use interactions, OSWorld (Xie et al., 2024) offers a unified environment for open-ended tasks involving various desktop and web applications. Specialized domains like marketing also feature benchmarks such as xbench (Chen et al., 2025f). Beyond specific domains, generalist agent benchmarks like AgentBench (Liu et al., 2023b), GAIA (Mialon et al., 2023), and TheAgentCompany (Xu et al., 2024b) evaluate broad problem-solving abilities across multiple knowledge domains and professional tasks, simulating real-world demands on general AI assistants.
@@
Beyond end-to-end task completion, assessing an agent’s underlying core competencies is crucial. These benchmarks evaluate fundamental capabilities that contribute to an agent’s overall intelligence and self-evolutionary potential. As for Planning, benchmarks such as PlanBench (Valmeekam et al., 2023), Natural Plan (Zheng et al., 2024), AutoPlanBench (Stein et al., 2025), and ACPBench (Kokel et al., 2025) comprehensively evaluate an agent’s ability to understand dynamic environments, devise strategies, decompose complex problems, and execute reasoning in various planning domains. For Tool Usage, simple benchmarks like ToolAlpaca (Tang et al., 2023) and ToolBench (Qin et al., 2023) test basic selection and parameter mapping, while more complex ones like ToolSandbox (Lu et al., 2024a), Seal-Tools (Wu et al., 2024), API-Bank (Li et al., 2023), T-Eval (Chen et al., 2023), $\tau$-Bench (Yao et al., 2024), AceBench (Chen et al., 2025a) simulate real-world scenarios involving multi-turn interactions, implicit state dependencies, and nested calls. Memory Management benchmarks such as LTMBenchmark (Castillo-Bolado et al., 2024), MemoryAgentBench (Hu et al., 2025), and StoryBench (Wan and Ma, 2025) evaluate the agent’s capacity to retain and utilize information across multi-turn interactions, dynamic scenarios, and long-range dependencies. For evaluating Multi-Agent Collaboration, benchmarks such as MultiAgentBench (Zhu et al., 2025) and SwarmBench (Ruan et al., 2025) assess coordination, communication, and emergent swarm intelligence in both collaborative and competitive settings.